# Sorting the Babble in Babel: Assessing the Performance of Language Detection Algorithms on the OpenAlex Database

### v3

@@ Abstract
This project aims to compare various language classification procedures, procedures combining various Python language detection algorithms and metadata-based corpora extracted from manually-annotated articles sampled from the OpenAlex database. Following an analysis of precision and recall performance for each algorithm, corpus, and language as well as of processing speeds recorded for each algorithm and corpus type, overall procedure performance at the database level was simulated using probabilistic confusion matrices for each algorithm, corpus, and language as well as a probabilistic model of relative article language frequencies for the whole OpenAlex database. Results show that procedure performance strongly depends on the importance given to each of the measures implemented: for contexts where precision is preferred, using the LangID algorithm on the greedy corpus gives the best results; however, for all cases where recall is considered at least slightly more important than precision or as soon as processing times are given any kind of consideration, the procedure combining the FastSpell algorithm and the Titles corpus outperforms all other alternatives. Given the lack of truly multilingual, large-scale bibliographic databases, it is hoped that these results help confirm and foster the unparalleled potential of the OpenAlex database for cross-linguistic, bibliometric-based research and analysis.
Keywords— Bibliographic Databases, Automatic Classification, Natural Language Processing
@@ 1 Introduction
At the crossroads of linguistics and computer science research, monolingual automatic language detection is nowadays considered a solved task in the literature, with state-of-the-art models achieving above human-level performance in as much as 1,366 languages (caswell2020language). Integrated in various real-world applications, language detection algorithms can also greatly enhance the organization, accessibility, and leveraging of information stored in databases. More precisely, by automatically identifying and categorizing the language of text attributes, these algorithms facilitate the sorting and indexing of multilingual data, which makes it easier to manage large, world-wide collections. Used in conjunction with machine translation algorithms, they can also support effective cross-language searches by allowing users to find pertinent information regardless of the language in which either the query or the relevant documents were written. Last but not least, language detection can benefit both data and resource management by respectively improving cross-linguistic data consistency and providing insights based on language distributions within the database.
Despite these advantages, the use of language detection algorithms in bibliographic databases remains as rare as scarcely documented. The main reason for this gap in implementation and research is both historical and infrastructural, as databases such as Scopus or the Web of Science have traditionally and quasi-exclusively focused on publications written in English (ammon2006language; archambault2005welcome; mongeon2016journal; demeter2020academic; vera2019web). This linguistic bias, while greatly reducing the need for multi- and cross-linguistic functionalities, also impacts the quality and scope of both research evaluation and research on research: being mostly blind to anything that eludes their monolingual lens, traditional bibliometric databases thus provide a fundamentally skewed view of the disciplinary and geographical landscape of the global scientific community (beigel2022batalla), which as a consequence distorts every initiative or investigation it informs and whose repercussions affect the neglected communities in various ways and degrees (tennant2020web; salatino2021fetichismo; finardi2022linguas; navarro2022rethinking).
Recent open-source alternative to the main traditional bibliometric databases, OpenAlex (OpAl) stands out from its predecessors by offering a broader landscape of scholarly communication (priem2022openalex). Built on an open codebase and from a variety of bibliometric projects, services, databases, and repositories, it has been shown to outperform both the Web of Science and Scopus in terms of document and journal frequency (alperin2024analysis; culbert2024reference; van2024oligopoly; jiao2023exclusively). Of particular relevance here, recent research has shown that OpAl also has a higher proportion of articles written in non-English languages than the two above-mentioned bibliographic databases (cespedes2025evaluating; vera2019web). In light of these considerations, not only does OpAl have the potential to offer a broader, multilingual, and as such more representative view of scholarly production at the global scale, but its stronger allophone content makes it also more susceptible to benefit from the functionalities and advantages language detection algorithms can offer to multilingual databases.
Regarding that last point, previous studies have however shown that language detection algorithm performance in multilingual context can vary significantly depending on the languages involved or on the specific characteristics of the dataset. In that perspective, effective multilingual database management requires linguistic classification procedures that are not only tailored to the specific linguistic features and metadata characteristics of the dataset, but also based on algorithms that are robust to noise in the data attributes considered (balazevic2016language; polasi2016combining; sujaini2024). In the specific case of OpenAlex, while a quality assessment of OpAl language metadata has already been carried out (cespedes2025evaluating), no attempt has yet been made yet to optimize the quality of its language metadata, that is, to find the best linguistic classification procedure for the articles it contains and indexes, based on its current metadata content and configuration.
Such comparative assessment of language detection procedures will be the focus of the present paper. More precisely, the objective here is to apply different algorithms on various article metadata corpora extracted from a manually-annotated sample of OpAl-indexed publications in order to find the algorithm-corpus combination that performs best over a selected set of evaluation measures. The following section gives a detailed account of the methodological outline designed for this project, followed respectively by an analysis and discussion of the results obtained.
@@ 2 Methods
Inferring the language of an indexed article based on its metadata is no simple procedure, as it involves two different computational tasks: detecting the language in which the metadata is written, and inferring the language of the article from that of its corresponding metadata. While closely related, these two tasks are however far from equivalent and can lead to opposite outcomes in certain contexts. For example, in the case of an article written in English but whose title is written in latin, an algorithm can only accurately predict the language of the article by failing to correctly identify the language of its title. Using more metadata attributes than just titles may help address such situations, but which set of metadata allows for optimal classification is most likely to change depending not only on the algorithm used, but also and perhaps more importantly on the language considered as well as on the performance measure considered; additionally, article metadata can be in different languages, which might substantially hinder algorithm performance.
In light of these considerations and in order to cover as much of the whole parameter space as possible, multiple linguistic classification procedures have been designed for this project. Following a series of rigorous data collection and processing operations, various metadata corpora were created by first collecting a set of manually- and linguistically-annotated samples to use as ground truth, then merging each article from the collected sample with various combinations of article metadata attributes considered relevant for language detection purposes, namely titles, abstracts, and journal names. Performance data for the resulting linguistic classifications is then collected by comparing predictions to observations for each annotated article in the sample. Procedure performance is then assessed at the language level first, using a series of evaluation measures commonly used in information retrieval, then at the database level, by simulating the overall performance of each algorithm using bayesian conjugate inferences over probabilistic confusion matrices. These different steps are detailed further in the following subsections.
@@ 2.1 Sample Collection
The ground truth for this study is based on the annotation work carried out by cespedes2025evaluating in their assessment of language metadata quality in OpAl. For that paper, a multilingual sample of indexed articles was manually annotated following a two step process. First, two samples of indexed articles were randomly drawn from the database, a first one containing 50 articles for each of the 55 languages present in the database (the resulting sample included only $2701$ articles, as some languages had less than 50 indexed articles) and a second containing 300 documents for each of the 11 most frequent languages in the database. Then, each article sampled was manually annotated by not only inspecting its title and abstract, but also by accessing the full text of the corresponding document, identify its language, and then determine if that paper indeed corresponds to a scientific paper; entries whose documents couldn’t be accessed were ignored (see (cespedes2025evaluating) for more information about these samples).
In addition to these two samples and in order to account for linguistically-unclassified articles indexed in OpAl, a sample of 300 such articles was collected for the purposes of the current project. Following extraction of this third subset, whose size relative to the previous two samples is proportional to the relative frequency of linguistically-unclassified articles in OpAl as of July 25th 2024, all articles thus sampled were annotated based on the same procedure as the one used for the two previous samples.
@@ 2.1.1 Corpus Building
As can be expected of a bibliographic database of such scale, OpAl contains articles from a variety of sources, from publishers and various scientific or professional organizations to conferences and self-archiving repositories. As a result, the type and quantity of metadata associated with articles indexed in the database can vary enormously: with regard to textual metadata in particular, some entries only include titles, whereas others include abstracts, journal names or both. Figure 1 shows the absolute and relative frequency distributions of all articles indexed in OpAl as of December 25th 2025, grouped by language and metadata configuration type. As regards to languages, the 11 most frequent languages in the database are directly represented in the figure, along with two additional categories which respectively include all articles classified in less frequent languages (other) and articles currently not classified linguistically (NA). As regards to configuration types, 4 different categories are distinguished: articles with titles, articles with titles and abstracts, articles with titles and journal names, and articles with titles, abstracts, and journal names. Absolute frequencies for languages and category types are respectively shown in millions (M) below the different y-axis and legend labels, while relative proportions for each metadata configuration type of each linguistic category are shown in percentages at the end of each corresponding horizontal bar. Finally, the x-axis was scaled algorithmically in order to better compare proportions and frequencies within and across language categories.
@@
As this figure shows, most articles in the database are written in English, as reported in cespedes2025evaluating, a majority of them also happen to include both abstracts and journal names as metadata. Following English, Asian languages form the most frequent linguistic group, followed by Romance languages. As for the different metadata attributes considered here, no clear cross-linguistic pattern can be found: while all articles have titles as attributes, metadata availability as regards to abstracts or journal names varies a lot across languages. Articles that include both their abstract and the name of the publishing journal represent a minority of indexed articles: besides English, only Chinese and Indonesian have a majority of indexed articles that include both articles and journal names. As for the remaining languages, most article entries only have journal names alongside titles, and a majority of Italian articles do not have any other relevant information other than titles.
For the purposes of the present project, such metadata variability is far from trivial. First of all, the performance of the algorithms used here should be expected to strongly correlate with input text length : the more text is given as input, the more accurate the predictions can be expected to be, regardless of the algorithm considered. Additionally, some language detection algorithms might be particularly sensitive to specific attribute features; given the constantly growing and changing nature of the OpAl database, algorithm performance at the database level at a given phase of its development could very well depend on the saliency and pervasiveness of these anti-patterns within the metadata configuration at any given time.
In light of these considerations and in order to find the language classification procedure whose predictive power is the most robust to such changes, different corpora were built for the sampled articles based on the metadata combination types used in figure 1.
Titles (T)
For each article, get the title.
Titles and Abstracts (A)
For each article, get its title and abstract, if available.
Titles and Journal Names(J)
For each article, get its title and the name of the journal it is published in, if available.
Greedy (G)
For each article, get its title, abstract, and the name of the journal it is published in, when either of the latter two is available.
Additional corpora were also generated based on all possible title, abstract, and journal name combinations; descriptions of these procedures and results obtained for each of them can be found in the appendix. Each of these corpora are built so as to include at least the title of every article in the annotated sample. This ensures that every procedure creates a complete corpus by assigning one document for each article sampled, thus allowing language detection algorithms to exhaustively partition the sample into the selected language categories. Finally, in order for each document to include complete sentences, text segments not ending with periods were punctuated accordingly.

### v4

@@ Abstract
This project aims to optimize the linguistic indexing of the OpenAlex database by comparing the performance of various Python-based language identification procedures on different metadata corpora extracted from a manually-annotated article sample  [^1] OpenAlex used the results presented in this article to inform the language metadata overhaul carried out as part of its recent Walden system launch [Demes_2025, Priem_2025].. The precision and recall performance of each algorithm, corpus, and language is first analyzed, followed by an assessment of processing speeds recorded for each algorithm and corpus type. These different performance measures are then simulated at the database level using probabilistic confusion matrices for each algorithm, corpus, and language, as well as a probabilistic modeling of relative article language frequencies for the whole OpenAlex database. Results show that procedure performance strongly depends on the importance given to each of the measures implemented: for contexts where precision is preferred, using the LangID algorithm on the greedy corpus gives the best results; however, for all cases where recall is considered at least slightly more important than precision or as soon as processing times are given any kind of consideration, the procedure that consists in the application of the FastText algorithm on the Titles corpus outperforms all other alternatives. Given the lack of truly multilingual large-scale bibliographic databases, it is hoped that these results help confirm and foster the unparalleled potential of the OpenAlex database for cross-linguistic and comprehensive measurement and evaluation.
keywords:
OpenAlex, Metadata Quality, Automatic Language Identification, Algorithm Performance Evaluation
@@ 1 Introduction
Automatic Language Identification (henceforth referred to as LI, as in [jauhiainen2024automatic]) generally refers to the task of predicting the natural language or languages present within a given written or speech sample [jauhiainen2019automatic, morris2020automated, jauhiainen2024automatic]. In the specific case of written samples, LI is closely related to the historical practice of deciphering, as both practices make extensive use of statistical techniques based on the compilation and comparison of linguistic patterns and frequencies. But most importantly, both tasks are formally solvable only insofar as their problem space is bounded: not only are the language samples they focus on generated by a grammar whose internal structure is finite, but more importantly, the set of candidate grammars that could have generated these language samples is itself limited [gold1967]. The extraordinary achievements of decipherment, from Egyptian hieroglyphs [champollion1822] to Linear B [ventris1952], were indeed achievable to the sole extent that decipherers operated within a bounded problem space: ancient scripts contain limited sign inventories, human languages draw from bounded morphological resources, and, crucially, the number of plausible languages an inscription could represent is itself finite, constrained by historical, geographical, and typological factors. In the case of automatic language detection, the problem space is even more constrained: the problem does not consist in retro-engineering long-lost grammars from language samples, but rather in identifying which grammar or grammars generated one or more input strings through comparisons within a closed set of pretrained language models. With a tight-bounded hypothesis space and the possibility of developing and relying on well-modeled distributional models, identifying the language or languages of written samples thus has the potential to become a relatively straightforward task from a statistical and computational perspective, at least in theory.
Attempts to automate this process were developed from the 1970s onward, focusing mainly on resource-rich, frequently-spoken languages [rau1974language, church1985, beesley1988, cavnar1994]. The global expansion of the internet has however significantly broadened the operational scope of LI, as algorithms are now often expected to identify any human language, including low-resource and endangered ones, both in real-time as well as in a wide variety of communication contexts, platforms, and formats [blodgett2022, joshi2020, solorio2014, wang2021]. Both performing and versatile, LI algorithms are now used in various tasks and for many different purposes. As multiple NLP tasks rely on models and tools that are tailored to specific languages, automatic Language Identification (LI) is often used as the first step of processing pipelines, from Machine Translation [goutte2014, zbib2012machine] to web-based, low-recourse language corpus building [jauhiainen2019automatic, jauhiainen2020building, jauhiainen2020experiments, caswell2020language]. However, and as the amount of recent literature on the subject demonstrates, LI is by no means a “solved task” [mcnamee2005language], since the complexities of real-world data still pose many challenges, particularly with respect to noise and multilingualism [jauhiainen2024automatic]. Perhaps more importantly, several potential areas that could benefit from LI, present new challenges, or offer opportunities for new developments are currently severely under-investigated. The case of multilingual information retrieval is particularly illustrative in this respect. Despite ongoing and significant advances in the field [zhang2023miracl, yang2024language, zhuang2023augmenting, elmahdy2024, goworek2025, ruder2017, jiang2020crosslingual, zhang2022mind], some obvious and crucial application areas remain surprisingly uninvestigated. The case of identifying the language of content of information records, that is, ”the language (or languages) used within the textual components of a resource (i.e. the language that a book is written in)” [morris2020automated, 1-2], certainly figures among them [knudson2014language, knudson2015automatic-key, chen2014investigation]. For example, a 2018 analysis of the British Library catalog revealed that 4.7 million library records lacked a language code and that only 30% records from the foundation catalogs were linguistically classified [morris2020automated]. Since very little data of this kind is available, it seems reasonable to assume that such inventories of collection metadata are seldom carried out, which could imply that the lack of information regarding the language of content of the British Library’s documents represents the norm rather than the exception.
A relatively similar situation can be observed in bibliometric databases, although for different reasons. For a long time, automatic language identification simply had little or no bibliometric relevance, as citation indices have traditionally focused on publications written in English [ammon2006language, archambault2005welcome, mongeon2016journal, demeter2020academic, vera2019web, vanleeuwen2001language, meneghini2007there]. From a language identification standpoint, this skewed language distribution made any attempt at optimizing language metadata rather pointless: one could simply but wrongly assume that all indexed articles are written in English and still get almost all of the cases right; in the same vein, any procedure aiming to adequately identify languages other than English would see their performance gain in that regard be outweighed by the necessarily higher number of English-written articles not identified as such. In a way, this situation can be likened to the one that prevailed within the field of Word Sense Disambiguation. For a long time, algorithms struggled to outperform the main sense heuristic (also called the most frequent sense heuristic), which simply consisted of assigning to each word its most frequent meaning [henrich2015word, abeysiriwardana2024survey, bevilacqua2021recent]. Of course, this situation was mainly due to the skewed nature of polysemy in natural language (the main sense of each word generally accounts for 70% to 90% of all use cases) [kilgarriff2004dominant, mccarthy2004finding, navigli2009word, arora2016detecting]), not to the contingent and biased nature of citation indexing practices. However, this more or less explicit partiality of traditional proprietary citation databases toward publications written in English does more than hinder the relevance and use of language identification algorithms. By only compiling references and citations of English-written publications, these databases provide a distorted disciplinary and geographical portrait of the research ecosystem, which misleads every initiative or investigation it informs and affects neglected communities in various ways and degrees [tennant2020web, salatino2021fetichismo, finardi2022linguas, navarro2022rethinking].
Within the context of this linguistic, cultural, and economic quasi-monopoly, the emergence and growth in popularity of the OpenAlex database feel truly transformational. Driven by its founding values of openness and comprehensiveness, this new citation index already offers the first truly broad and multilingual picture of the global scholarly landscape [priem2022openalex]. Based on an open codebase and on a variety of bibliometric projects, services, databases, and repositories, OpenAlex has been shown to outperform both the Web of Science and Scopus in terms of document and journal frequency [alperin2024analysis, culbert2024reference, van2024oligopoly, jiao2023exclusively]. Of particular relevance here, recent research has shown that OpenAlex also has a higher proportion of articles written in non-English languages than all other bibliometric databases [cespedes2025evaluating, vera2019web]. This stronger allophone content, while making OpenAlex the first citation index likely to benefit from the functionalities offered by automatic language identification, also opens the door for a first thorough evaluation of LI performance in bibliometric context. Such comparative assessment of LI procedures is the focus of the present paper. More precisely, the objective here is to apply different algorithms on various article metadata corpuses extracted from a manually annotated sample of OpenAlex-indexed publications in order to find the algorithm-corpus combination that performs best over a selected set of evaluation measures. The next section will give an overview of the unique challenges posed by Open Alex to current-day LI research, followed by a descriptive and justificatory account of the methodological apparatus designed for the project. The project’s results are then presented, analyzed, and discussed in the last two sections.
2 Research Context
Both in structure and content, OpenAlex however has the potential to pose some serious challenges to LI research. First of all, not all text attributes that can be found in the article metadata carry language signals, that is, observable cues in text that reflect the grammar, vocabulary, and orthography of of a given language and makes its identification possible [elliott2000language]. In the present project, three different metadata attributes have been considered to carry adequate language signals and thus included in the analysis: title, abstracts, and journal names. However, all three metadata attributes are not always present in all records; some entries only include titles, whereas others include abstracts, journal names, or both. These completeness issues could undoubtedly affect the performance of algorithms whose robustness fails to cope with such metadata irregularities. Also, titles and journal names often contain only a handful of words, thus making them poor carriers of linguistic signal and thus difficult targets for even the best models [jauhiainen2019automatic]. Moreover, research has shown that the severity of this language signal size problem increases in proportion to the number of languages to identify [baldwin2010language], which makes the task of identifying the language of content of indexed books and periodicals all the more problematic [morris2020automated], especially in a database as linguistically diverse as OpenAlex [cespedes2025evaluating]
Perhaps even more problematic, OpenAlex is replete with mixed-language content, one of the most challenging issues current-day LI algorithms aim to address. Theoretically-grounded in sociolinguistics, current LI research draws a distinction between two broad types of multilingual context: code-switching, which refers to the alternation between languages across sentence boundaries, and code-mixing, which refers situations in which elements from more than one language are combined within a single sentence, from simple loanwords to linguistically hybrid, yet grammatically coherent crosslinguistic intertwinings [muysken2000bilingual]. Code-mixing is generally considered more complex than its counterpart: cognates (words in different languages that have the same origin and often similar forms), interlingual homographs (words that are spelled the same in two or more languages, but are not etymologically or semantically related), loanwords (words borrowed from another language), and internationalisms (words used globally in many languages) can all undermine the performance of LI algorithms. Such code-mixing can prove especially challenging in the context of scholarly metadata, as it is perhaps more pregnant than other linguistic contexts: proper nouns, technical terminology, words with greek or latin roots (et al., in vitro, quantum,…), and acronyms (CRISPR, AI, LASER, RADAR, SONAR,…), which are often integrated as such across languages, tend to dominate titles, abstracts, and journal names. In addition, modern metadata often relies on publisher-supplied English glosses in parentheses or Latin transliterations of person names, geographic names, or other content not originally written in Latin characters, all which further ambiguate language signals. LI is extremely difficult in such contexts, as language cues point in multiple directions [winata2023decades, goswami2020udldi, morris2020automated]. Moreover, the shorter the number of characters, the more ambiguous or weaker the linguistic cue, and thus the harder to adequately identify the language.
As for code-switching, records frequently include titles followed by a translation in parentheses, or abstracts in two or more languages following each other in sequence. In such cases, boundaries exist between sentences, but metadata schemas tend to not tag them explicitly and include them in a single attribute. A different sort of code-switching can also be found in mixed-language attribute databases like OpenAlex: not only does the same attribute can be in multiple languages across the database, but a single record’s attributes can also be in different languages. This happens for example when a journal’s name or an article title is in a different language than the language of content of the article; Latin and Greek are commonly used for this purpose [morris2020automated], but modern languages can also be used for that very specific purpose.
For LI research in general, the presence of code-mixing and code-switching implies that the standard “one language per document” assumption, more or less explicitly assumed by most existing research in the field, can no longer be reasonably maintained, which means that systems must shift from monolingual, record-level modeling to finer-grained representations able to handle intra-record multilinguality. This requires rethinking database schemas, information retrieval architectures, and NLP pipelines to support heterogeneous language attributes, dynamic language-specific processing, and multi-label evaluation across all phases of database development and maintenance [jauhiainen2019automatic, Jauhiainen2024PlenarySessions, zhang2023miracl]. OpenAlex does not require such a database overhaul, to the extent that the objective is the same as that of the British Library catalogue project: to identify the main language of content of indexed publications [morris2020automated]. However, this monolingual purpose can still be considerably affected by the mixed linguistic content of a substantial part of OpenAlex’s records. And here, as in other areas that rely on LI, failure to adequately handle multilingualism will unavoidably result in over-representation of dominant languages and lead to serious discoverability problems, since all records whose content language is not properly identified will be omitted from any language-specific user search or query [morris2020automated]. In the context of the predominance of English as the lingua franca of Science, both shortcomings will thus tend to reinforce the structural dominance of English in global scholarly productivity and impact [visser2021large, morris2020automated], henceforth reinforcing global scholarly production inequalities and distorting attempts to measure international research diversity or linguistic coverage of databases [visser2021large]. In light of this, the use of LI algorithms on a multilingual bibliometric database like OpenAlex is more than a simple preprocessing step, but rather a crucial component of responsible bibliometric methodology and epistemic justice.
A final and broader challenge for the application of LI on a large database like OpenAlex pertains to algorithmic feasibility, scalability, incrementability, and implementability, requirements of crucial importance to large-scale databases in general. As OpenAlex contains hundreds of millions of records, the spatial and temporal complexity of LI algorithms, that is, how memory usage and execution time grow as input size increases, is at least as important as how well they manage to identify the language of content of every publication indexed in the database. Moreover and as is often the case with real-world, multilingual systems in general, LI procedures need to be re-run incrementally on the whole database, as its language metadata keeps getting corrected or enriched through the integration of new attributes and records. While increasing the criticality of spatio-temporal constraints, this constant need for updates also makes unusable any procedure based on the retraining of large models.
Considered altogether, the different challenges discussed in this section provide a clear but sobering picture of the situation: OpenAlex record attributes are often short, noisy, inconsistently labeled, with language shifts often occurring within a single attribute, between different attributes of the same record, and even within the same attribute across different records. Given the extensive of bibliometric databases for high-stakes analytics, identifying the language of content of indexed articles is not only an extremely delicate and difficult task, but also one that has the potential to severely distort any picture of the scientific landscape offered by the database, which can in turn profoundly affect science mapping as well as research measurement and evaluation. Ensuring the quality of the linguistic metadata in the OpenAlex database, for the challenges it presents as well as for the issues that arise from it, thus represents a unique and crucial opportunity from both a bibliometric and natural language processing perspective. The next section will describe the method designed here to achieve this goal.
@@ 3 Methods
Inferring the language of content of an indexed article based on its metadata is no simple procedure, as that language has to be indirectly inferred from the language signals contained in its metadata. Using more metadata attributes than simply titles may help in this task, but the problem of finding the set of attributes that allows for optimal classification is most likely to change depending not only on the algorithm used, but also and perhaps more importantly on the language considered as well as on the performance measure used. In light of these considerations and to cover as much of the whole parameter space as possible, multiple linguistic classification procedures have been designed for this project. A linguistic classification procedure is here defined as the unique combination of an automatic language identification algorithm and a corpus type.
Following various data collection and processing operations, the different metadata corpus types used in this project were generated by first collecting a set of manually and linguistically annotated samples to use as ground truth, then merging each article from the collected sample with different corpora representing all possible combinations of article metadata attributes considered relevant for LI purposes, namely titles, abstracts, and journal names  [^2] Author- and journal-supplied keywords were not indexed in the first version of OpenAlex, most probably for copyright reasons [courtney2024no, carroll2009copyright]. Beyond these legal considerations, however, it could be argued that the inconsistent indexing of keywords, along with their metalinguistic as well as linguistically unnatural and ambiguous nature, all constitute strong arguments against their use for article language identification purposes.
The different language classification procedures evaluated in this project were then completed by applying each algorithm selected to each and every metadata corpus generated. We then collected performance data for the resulting linguistic classifications by comparing predictions with observations for each annotated article in the sample. Procedure performance is then assessed at the language level first, using a series of evaluation measures commonly used in information retrieval, then at the database level, by simulating the overall performance of each algorithm under different indicator weighing regimes using Bayesian conjugate inferences over probabilistic confusion matrices. These different steps are detailed further in the following subsections.
@@ 3.1 Sample Collection
The ground truth for this study is based on the annotation work performed by cespedes2025evaluating in their assessment of the quality of language metadata in OpenAlex. For that paper, a multilingual sample of indexed articles was manually annotated following a two-step process. First, two samples of indexed articles were randomly drawn from the database, a first one containing 50 articles for each of the 55 languages present in the database (the resulting sample included only $2701$ articles, as some languages had less than 50 indexed articles) and a second containing 300 documents for each of the 11 most frequent languages in the database. Then, each article sampled was manually annotated by not only inspecting its title and abstract, but also by accessing the full text of the corresponding document, identify its language, and then determine if that paper indeed corresponds to a scientific paper; entries whose documents couldn’t be accessed were ignored (see [cespedes2025evaluating] for more information about these samples).
In addition to these two samples and in order to account for linguistically-unclassified articles indexed in OpenAlex, a sample of 300 such articles was collected for the purposes of the current project. Following extraction of this third subset, whose size relative to the previous two samples is proportional to the relative frequency of linguistically-unclassified articles in OpenAlex as of July 25th 2024, all articles thus sampled were annotated based on the same procedure as the one used for the two previous samples.
@@ 3.1.1 Corpus Building
As can be expected of a bibliographic database of such scale, OpenAlex contains articles from a variety of sources, from publishers and various scientific or professional organizations to conferences and self-archiving repositories. As a result, the type and quantity of metadata associated with articles indexed in the database can vary enormously: with regard to textual metadata in particular, some entries only include titles, whereas others include abstracts, journal names, or both. Figure 1 shows the absolute and relative frequency distributions of all articles indexed in OpenAlex as of December 25th, 2025, grouped by language and metadata configuration type. As regards to languages, the 11 most frequent languages in the database are directly represented in the figure, along with two additional categories, which, respectively, include all articles classified in less frequent languages (other) and articles currently not classified linguistically (NA). Regarding configuration types, four different categories are distinguished: articles with titles, articles with titles and abstracts, articles with titles and journal names, and articles with titles, abstracts, and journal names. Absolute frequencies for languages and category types are respectively shown in millions (M) below the different y-axis and legend labels, while relative proportions for each metadata configuration type of each linguistic category are shown in percentages at the end of each corresponding horizontal bar. Finally, the x-axis was scaled algorithmically in order to better compare proportions and frequencies within and across language categories.
@@
As this figure shows, most of the articles in the database are written in English, as previously reported in [cespedes2025evaluating], and a majority of these English articles include both abstracts and journal names as metadata. Asian languages form the second most frequent linguistic group, followed by Romance languages. As for the different metadata attributes considered here, no clear cross-linguistic pattern can be found: while all articles have titles as attributes, metadata availability as regards to abstracts or journal names varies a lot across languages. Articles that include both their abstract and the name of the publishing journal represent a minority of indexed articles: besides English, only Chinese and Indonesian have a majority of indexed articles that include both articles and journal names. As for the remaining languages, most article entries only have journal names alongside titles, and a majority of Italian articles do not have any relevant information other than titles.
For the purposes of the present project, this metadata variability is far from trivial. First of all, the performance of the algorithms used here should be expected to strongly correlate with input text length: the more text is given as input, the more accurate the predictions can be expected to be, regardless of the algorithm considered. Additionally, some LI algorithms might be particularly sensitive to specific attribute features; given the constantly growing and changing nature of the OpenAlex database, algorithm performance at the database level at a given phase of its development could very well depend on the saliency and pervasiveness of these antipatterns within the metadata configuration at any given time.
In light of these considerations and to find the language classification procedure whose predictive power is most robust to such changes, different corpora were built for the sampled articles based on the metadata combination types used in figure 1.
Corpus Type Procedure
Titles (T) For each article, get the title.
Titles & Abstracts (A) For each article, get its title; if available, get its abstract.
Titles & Journal Names (J) For each article, get its title; if available, get the name of the journal.
Greedy (G) For each article, get its title; if available, get its abstract and the name of the journal.
Table 1: Description of the corpus types
Each of these corpora is built so as to include at least the title of every article in the annotated sample. This ensures that every procedure creates a complete corpus by assigning one document to each article sampled, which allows LI algorithms to exhaustively partition the sample into the selected language categories. Finally, in order to syntactically separate document attributes in each corpus, text attributes not ending with periods were punctuated accordingly.